An exponentially growing culture has a specific growth rate constant, K = 2.4 day-1. What is...

  1. An exponentially growing culture has a specific growth rate constant, K = 2.4 day-1. What is the generation time? A test tube of fresh medium is inoculated with initial cell population of 6 x 105 cells/mL. Assuming no lag time, how long will it take for the culture to reach 8 x 109 cells/mL.
Answer #1

Mean generation time = 1 / K

= 1 / 2.4

= 0.4166 day

No= 6 x 105 cells/mL

Nt = 8 x 109 cells/mL

n = Number of generations

So n = ( log Nt - log No) / log 2

= (9.903 - 5.778) / 0.301

= 4.125 / 0.301

= 13.704 generations

Hence t = n / K

= 13.704 / 2.4

= 5.71 days ( ~ 5 days 17 hrs)